OpenStudy (anonymous):

Add parentheses to the equation below so that x- -5 5x4-6/2-x (/ are division =) A.(5x4)-6/2=x B. 5x(4-6/2)=x C. 5x4-(6/2)=x D. 5x(4-6)/2=x PLEASE HELP!

OpenStudy (anonymous):

is that 5x^4 or 5 multiplied by 4 ?

OpenStudy (anonymous):

5 x 4 - 6 / 2 - x @kelliegirl33

OpenStudy (anonymous):

wait I see hold on

OpenStudy (anonymous):

so the outcome will be x = - 5 ?

OpenStudy (anonymous):

5 * 4 - 6 / 2 - x (this * will be used a multiplication symbol) @kelliegirl33

OpenStudy (anonymous):

(a) (5 x 4) - 6/2 = x 20 - 6/2 = x 40/2 - 6/2 = x 34/2 = x 17 = x Not A (b) 5 (4 - 6/2) = x 5( 8/2 - 6/2) = x 5( 2/2) = x 10/5 = x 2 = x Not B (c) 5 x 4 - (6/2) 20 - 6/2 this is the same as A. Not C (d) 5 (4 - 6)/2 = x 5( - 2)/2 = x 5 (-1) = x - 5 = x BINGO.....it is D

OpenStudy (anonymous):

Thank you so much for your time a effort! :) @kelliegirl33

OpenStudy (anonymous):

no problem.....glad to help :)

OpenStudy (anonymous):

I'm your fan may I come to you for any more questions?

OpenStudy (anonymous):

yes you can....I can't promise that I will know the answer...but I will try :)

OpenStudy (anonymous):

so your sure this ones right xD

OpenStudy (anonymous):

yes :) I worked them out......if you don't trust my answer, get a second opinion. This is your grade, not mine, so you want to be sure.

OpenStudy (anonymous):

I trust you girl!! your my go to girl now :)